OpenStudy (anonymous):

Solve for x 1/x+2/z=2/y

OpenStudy (anonymous):

x=(2/y-2/z)^-1

OpenStudy (anonymous):

How does the power to negative one happen?

OpenStudy (cruffo):

Yah, that would work, but probably not the answer you were expecting :)

OpenStudy (cruffo):

Here's the idea: Step 1: isolate the term with the x in it: \[\frac{1}{x} = \frac{2}{y} - \frac{2}{z}\]

OpenStudy (anonymous):

multiply both by x/1?

OpenStudy (cruffo):

Yes, you could continue that way.

OpenStudy (cruffo):

But that will create a bit of a mess.

OpenStudy (anonymous):

ok, how should i continue?

OpenStudy (cruffo):

Consider this, if we had \[\frac{1}{x} = \frac{2}{3}\] Then we could say that \[\frac{x}{1} = \frac{3}{2}\]

OpenStudy (cruffo):

Flip both fractions :)

OpenStudy (cruffo):

Suppose you add the two fractions on the right hand side.

OpenStudy (anonymous):

2/z and 2/y?

OpenStudy (anonymous):

after flipping everything though?

OpenStudy (cruffo):

\[\frac{2}{y} - \frac{2}{z} = \frac{2z - 2y}{yz}\]

OpenStudy (cruffo):

You want to add first, then flip.

OpenStudy (anonymous):

I can add two different fractions together?

OpenStudy (anonymous):

so, 1/x + 2/z = 2/y i should first subtract 2/z so i get 1/x = 2/y - 2/z

OpenStudy (cruffo):

Trick only works when you have fraction = fraction. Not a sum or difference. To add fractions with different denominators you need a common denominator. In this case just multiply the two denominators together

OpenStudy (anonymous):

so I cross multiply the isolated fractions to get that crazy one? and have x/1 = 2z-2y/yz

OpenStudy (cruffo):

Good. Now you can flip!

OpenStudy (cruffo):

Hum. in the previous, you had the left already flipped!

OpenStudy (anonymous):

sorry it was supposed to be 1/x= 2z-2y/yz

OpenStudy (cruffo):

Yah, ok. Good!

OpenStudy (anonymous):

then I go x/1 = yz/2z-2y

OpenStudy (cruffo):

You got it!

OpenStudy (anonymous):

ok...so how do I get x alone? sorry! Im so confused

OpenStudy (cruffo):

x/1 is just x :)

OpenStudy (anonymous):

ok

OpenStudy (cruffo):

If you want, you could follow your original train of thought: After you isolate the term with the x in it (step 1), you could multiply every term on both sides by x. But then you will still have to add the fractions (with more letters in them!!!). After that would would have to factor!

OpenStudy (anonymous):

yeah, that would be harder.

OpenStudy (cruffo):

Cheers!